Give an example of a function of two variables such thatf(0,0) = 0 butfis NOT continuousat (0,0). Explain why the functionfis NOT continuous at (0,0).


1
Expert's answer
2021-12-15T05:09:03-0500

Consider the function "f:\R\times\R\to\R,\ f(x,y)=\begin{cases}0, \text { if }(x,y)=(0,0)\\ 1, \text { if }(x,y)\ne(0,0)\end{cases}."

Taking into account that "\lim\limits_{x\to 0,\\ y\to 0}f(x,y)=1\ne 0=f(0,0)," we conclude that the function is not continuous at "(0,0)."